# Automatic infant 2D pose estimation from videos: comparing seven deep neural network methods

### v2

@@ Acknowledgements
In addition to the real infants annotated data mentioned in the main manuscript, additional images were annotated.
We manually annotated 720 additional images from eight additional videos of the same two infants. The images were selected in a similar manner as described in the main manuscript, but were processed only with video input by the methods, hence why they are not included in the main manuscript where the results between videos and images input are compared. Thus, we reach a total of 1440 annotated images from 16 videos (191 862 images) that were processed as video inputs by all methods.
We also annotated 900 additional consecutive images from a single video, the one identified as "AA_17w" in several tables in the Supplementary Materials. This sequence was selected due to the presence of typical hand movements and self-touch occurrences that we are interested in studying in the future. In retrospective, we observed that this video is the one where all the methods performed the best (except for DeepLabCut and MediaPipe, see Tab. ST. 4), most likely because the camera is properly positioned and angled above the infant (see Figure 1., b) in the main manuscript), and the infant does not perform many complex leg movements. As such, it could be considered to have optimal conditions and show the upper bound performance of the methods on infants in supine position. Due to how this sequence was arbitrarily chosen and because it only concerns a single infant at a single age, we decided to not include it in the main manuscript, and leave it as an extra in the Supplementary Materials.
@@ Results
Settings for detection selection with redundant detections
In most applications, when there are redundant detections, a selection process is necessary to choose one of them for further processing. This can only be done by using information that is provided by the pose estimation methods alongside the keypoints, such as the rank order in which the detections are output, or their scores and confidences. We computed Euclidean distances between estimations and their ground truth, the first processing step at the base of the OKS and Neck-MidHip computations, under each of these three detection selection settings:
  • Det 1: using the first ranked detection provided by the method. Redundant detections proposed by the method are ignored.
  • Det 2: the average Euclidean distance is computed for each detection of a given image. Then, the detection with the shortest distance is selected. This corresponds to the best detection that the method can offer, though it is not available without ground truth.
  • Det 3: using the detection with the highest score, ignoring the rank order. For HRNet TD that did not provide scores, we used the score of the bounding-box provided by its detector as a substitute. For ViTPose, we used the median of its confidences.
These settings are used to 1) verify the general assumption that the detection with the highest score is always ranked first, and if not, 2) which of the two available selection metrics (rank or score) matches the best detection proposed by the methods and is a better choice.
Minimal differences were observed when comparing the results for all relevant metrics between the three detection selection settings, namely, the first-rank detection, the optimal detection, and the highest-scored detection. DeepLabCut, MediaPipe and OpenPose yielded identical results, as they only provided one detection. AlphaPose and ViTPose showed a few cases where the first-ranked detection did not have the highest score, though for ViTPose it might be due to how we estimated its score as the median of the confidences: a different estimation might have led to different or no changes. Both first-ranked and highest-score detections had differences with the optimal detection, meaning that there is no guaranteed way to select the best available detection for any of the methods when no ground-truth is available. Generally, using the highest-scored detection resulted in the closest performance to the optimal detection for all methods. Hence, the whole manuscript focuses on showing the results from the highest-scored detections only.
Results summary for the dataset of 1440 real infants annotated images processed by video input
The summary of the results including all 1440 annotated images from the 16 recordings is described in the Tab. ST 2.
We observe similar results as in the main manuscript, with a tendency for slightly higher AP, AR and OKS, and lower missing data, but also lower correlations between score and OKS values.
Metrics AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
OKS 0.88$\pm$0.07 0.11$\pm$0.12 0.87$\pm$0.09 0.32$\pm$0.20 0.90$\pm$0.06 0.92$\pm$0.05 0.81$\pm$0.17 0.92$\pm$0.04
AP 66.5 0.8 25.3 0.6 77.8 59.1 59.9 86.7
AR 76.9 0.3 79.1 4.0 85.9 88.5 66.8 90.0
Missing data 3.3% 0% 0% 5.6% 0% 0.9% 6.1% 0.1%
Redun. det. 4.6% 0% 259.5% 0% 5.1% 54.0% 0% 19.7%
Sco.-OKS corr. 0.26 0.08 0.21 N/A 0.51 0.47 0.35 0.39
Table 2: Summary of the highest-scored detection results for the full set of 1440 real infants annotated images processed by video input. For score-OKS Spearman Rank Coefficient Correlations, all p-values < 0.005.
Results summary for the 900 continuous annotated images from AA_17w
The results summary for the 900 continuous manually-annotated images from a single recording session, AA_17w, are summarized in Supplementary Tab. ST. 3.
The AP, AR, and OKS values are higher than the averages found in the main manuscript or in Supplementary Tab. ST 2, while the average Neck-MidHip errors are lower, except for DeepLabCut and MediaPipe.
Concerning the correlations, we observe that they are particularly low, especially for Detectron2, where they are negative.
Input Metrics AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
Images OKS 0.90$\pm$0.04 N/A 0.93$\pm$0.04 0.40$\pm$0.10 0.94$\pm$0.03 0.95$\pm$0.03 0.93$\pm$0.05 N/A
AP 77.9 N/A 81.4 0.1 91.2 66.4 87.8 N/A
AR 81.7 N/A 90.0 2.4 94.3 95.2 90.6 N/A
Neck-MidHip error 5.0$\pm$1.7N/A 4.7$\pm$2.424.7$\pm$12.04.0$\pm$2.03.9$\pm$2.24.5$\pm$1.9N/A
Sco.-OKS corr. 0.11 N/A -0.35 N/A 0.04 ($p=0.26$) 0.30 0.06 ($p=0.06$) N/A
Videos OKS 0.90$\pm$0.04 0.08$\pm$0.06 0.92$\pm$0.04 0.40$\pm$0.10 0.94$\pm$0.03 0.95$\pm$0.03 0.93$\pm$0.05 0.95$\pm$0.02
AP 76.9 0.0 34.2 0.3 90.8 63.9 87.3 91.8
AR 81.1 0.0 89.9 2.5 93.8 95.1 90.2 94.5
Neck-MidHip error 5.1$\pm$1.879.3$\pm$15.34.9$\pm$2.524.8$\pm$0.124.0$\pm$2.03.9$\pm$2.24.6$\pm$1.93.9$\pm$2.0%
Sco.-OKS corr. 0.10 ($p<0.006$) 0.11 -0.32 N/A 0.01 ($p=0.70$) 0.28 0.10 0.18
Table 3: Summary of the highest-scored detections results from each metric for the set of 900 consecutive annotated images from video AA_17w. For score-OKS Spearman Rank Coefficient Correlations, all p-values < 0.005 when not stated.
Object Keypoint Similarity (OKS)
The average OKS of individual videos can be seen in Tables ST 4 and ST 5. It can be used to identify the most challenging or easiest videos for each method.
As the synthetic infants from the MINI-RGBD dataset come with an estimation of the complexity of their sequence, we can observe that the performance between the easy group (IDs 1-4) and the medium group (5-9) seems close, except for synthetic infant 1 that some methods seem to struggle with, and synthetic infant 9 for which all methods show a drop of performance. Synthetic infants 7 and 8 are handled differently by the methods, some keeping their performance levels, while others display drops of performance. However, for all methods, we can observe a performance drop between the easy and medium video group and the difficult video group (IDs 10-12).
Real Video ID AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
Images AA_8w 0.86 N/A 0.78 0.50 0.81 0.89 0.79 N/A
AA_17w 0.90 N/A 0.92 0.39 0.94 0.94 0.92 N/A
TH_8w_st1 0.89 N/A 0.88 0.43 0.90 0.91 0.88 N/A
TH_8w_st2 0.90 N/A 0.89 0.39 0.91 0.93 0.89 N/A
TH_8w_st3 0.88 N/A 0.89 0.32 0.90 0.91 0.89 N/A
TH_15w 0.77 N/A 0.86 0.17 0.90 0.92 0.80 N/A
TH_19w 0.87 N/A 0.86 0.19 0.90 0.93 0.88 N/A
TH_25w 0.89 N/A 0.87 0.50 0.93 0.93 0.88 N/A
Mean 0.87 N/A 0.87 0.39 0.90 0.92 0.87 N/A
Videos AA_8w 0.86 0.31 0.77 0.52 0.81 0.89 0.79 0.94
AA_11w 0.86 0.19 0.85 0.19 0.90 0.91 0.85 0.91
AA_13w 0.90 0.48 0.91 0.19 0.91 0.94 0.81 0.94
AA_17w 0.90 0.09 0.92 0.37 0.94 0.94 0.92 0.95
AA_19w 0.89 0.02 0.90 0.15 0.93 0.93 0.91 0.93
TH_8w_st1 0.88 0.15 0.87 0.42 0.90 0.91 0.87 0.91
TH_8w_st2 0.90 0.07 0.88 0.39 0.91 0.93 0.90 0.93
TH_8w_st3 0.88 0.10 0.89 0.36 0.90 0.90 0.89 0.92
TH_10w_st1 0.90 0.15 0.89 0.47 0.91 0.93 0.89 0.93
TH_10w_st2 0.84 0.08 0.87 0.33 0.90 0.89 0.80 0.90
TH_12w 0.89 0.20 0.88 0.40 0.92 0.90 0.84 0.93
TH_15w 0.75 0.04 0.86 0.19 0.90 0.92 0.58 0.92
TH_17w 0.87 0.04 0.86 0.07 0.90 0.88 0.83 0.89
TH_19w 0.88 0.05 0.86 0.11 0.90 0.92 0.52 0.93
TH_23w 0.87 0.06 0.82 0.24 0.91 0.91 0.68 0.93
TH_25w 0.88 0.17 0.87 0.49 0.93 0.93 0.76 0.94
Mean 0.88 0.11 0.87 0.32 0.90 0.92 0.81 0.92
Table 4: Average OKS values over all manually-annotated images for each method and input type on real infants.
Synth. Video ID AlphaPose DeepLabCut Detectron 2 MediaPipe HRNet BU HRNet TD OpenPose ViTPose
Images Syn. 1 0.83 N/A 0.84 0.34 0.61 0.90 0.84 N/A
Syn. 2 0.87 N/A 0.83 N/A 0.90 0.91 0.86 N/A
Syn. 3 0.89 N/A 0.90 0.52 0.90 0.91 0.89 N/A
Syn. 4 0.90 N/A 0.90 0.62 0.90 0.90 0.84 N/A
Syn. 5 0.89 N/A 0.87 0.59 0.90 0.90 0.85 N/A
Syn. 6 0.86 N/A 0.88 0.57 0.91 0.90 0.88 N/A
Syn. 7 0.89 N/A 0.89 0.59 0.91 0.91 0.88 N/A
Syn. 8 0.85 N/A 0.85 0.45 0.87 0.89 0.84 N/A
Syn. 9 0.73 N/A 0.82 0.52 0.86 0.90 0.72 N/A
Syn. 10 0.79 N/A 0.79 0.39 0.76 0.81 0.75 N/A
Syn. 11 0.84 N/A 0.87 0.55 0.87 0.88 0.86 N/A
Syn. 12 0.67 N/A 0.66 0.35 0.75 0.78 0.75 N/A
Mean 0.84 N/A 0.84 0.50 0.83 0.88 0.83 N/A
Videos Syn. 1 0.76 0.37 0.85 0.22 0.59 0.87 0.85 0.89
Syn. 2 0.86 0.46 0.82 N/A 0.89 0.90 0.84 0.89
Syn. 3 0.88 0.49 0.88 0.51 0.88 0.89 0.87 0.89
Syn. 4 0.86 0.40 0.86 0.61 0.82 0.84 0.80 0.88
Syn. 5 0.88 0.59 0.85 0.57 0.87 0.90 0.85 0.90
Syn. 6 0.86 0.51 0.87 0.57 0.90 0.90 0.86 0.89
Syn. 7 0.87 0.55 0.87 0.58 0.89 0.89 0.86 0.88
Syn. 8 0.83 0.36 0.83 0.44 0.85 0.87 0.82 0.85
Syn. 9 0.48 0.38 0.74 0.42 0.78 0.87 0.70 0.89
Syn. 10 0.78 0.35 0.75 0.38 0.74 0.78 0.72 0.83
Syn. 11 0.83 0.50 0.85 0.54 0.86 0.86 0.84 0.84
Syn. 12 0.56 0.19 0.59 0.35 0.69 0.75 0.69 0.79
Mean 0.81 0.43 0.81 0.47 0.81 0.86 0.81 0.87
Table 5: Average OKS values over all manually-annotated images for each method and input type on synthetic infants.
@@
Figures showing the average Neck-MidHip errors, including DeepLabCut, for each individual keypoint with available ground truth are shown in Fig. SF 1.
Figure 1: Average Neck-MidHip errors for each keypoint with available ground truth. The centre of the circles is the ground-truth position for that keypoint. The radius of each circle shows the average amplitude of the errors, scaled to the Neck-MidHip segment. The colors separately represent each pose estimation method. (a) Real infants, video input (720 annotations); (b) Real infants, video input (1440 annotations); (c) Synthetic infants, video input
The details of the overall average Neck-MidHip errors across all keypoints with their standard deviations for each method are shown in Tab. ST. 6. The best method is ViTPose, achieving average errors as low as 6.0$\pm$2.7of the Neck-MidHip segment. HRNet Top-Down is slightly behind. However, the variability between each keypoint is generally high: the standard deviations across all keypoints are often above one third, and sometimes even half, of the average error. This can also be see more visually in the main manuscripts’ Fig. 2 and in the Supplementary Materials Fig. SF. 1.
@@
Real Images [720] $8.3\pm 3.1$ N/A $10.2\pm 5.2$ 38.0$\pm$13.8 $7.8\pm 3.8$ 6.7$\pm$3.6 $8.8\pm 4.1$ N/A
Videos [720] $8.3\pm 2.9$ $99.6\pm 25.9$ $10.5\pm 5.4$ 35.1$\pm$13.7 $7.9\pm 3.9$ $6.7\pm 3.6$ $12.3\pm 4.3$ 6.0$\pm$2.7
Videos [1440] $7.6\pm 2.8$ $103.6\pm 20.5$ $9.4\pm 4.3$ 44.6$\pm$15.5 $7.1\pm 3.4$ $6.6\pm 3.5$ $10.7\pm 3.4$ 6.0$\pm$2.9

### v3

@@ Methods
Unlike for adult pose estimation, datasets of infant recordings with annotated keypoints are scarce and usually not publicly shared. For this study, we used the only two available public datasets that we know of, as well as two datasets from our own video recordings, which we manually annotated (occluded keypoints were not annotated).
Our main comparison datasets concern infants in supine position, with a single infant in every image.
  • Real infants supine dataset. 720 annotated images (90 per video) from 8 videos (59200 images total) of two infants, followed longitudinally between 2 and 6 months of age. The annotations were made using the DeepLabCut labeling tool. 480 images (60 per video) were chosen by taking the first ten images every 100 images, from 0 to 9, then 100 to 109, until 509. The remaining 240 images (30 per video) were selected through DeepLabCut’s integrated selection tool using k-means clustering to pick the least similar images in each video. Fourteen keypoints were annotated, corresponding to the COCO keypoints minus the ears and the nose. Our dataset is entirely comprised of infants recorded at home by their parents, each video being different from the others in terms of camera angles, lighting conditions, as well as background and clothes colors and type.
  • Synthetic infants – MINI-RGBD. Based on recordings of real infants, Hesse and colleagues have trained a model capturing infant shape and posture distributions which they subsequently used to produce and render synthetic sequences of artificial infants under 7 months in supine position—the MINI-RGBD dataset [54]. There are 12 synthetic infants with 1000 images each. Ground truth for 25 keypoints is available. Of these 25 keypoints, 13 are in common with COCO’s 17 keypoints and were used to evaluate the methods. Each of the twelve synthetic sequences is unique, including backgrounds with large variations between sequences (see Fig. 1 for two examples, or see the original paper [54]).
In addition, we used two datasets of annotated images of infants in different and more complex contexts:
  • Real infants on the lap dataset. 359 annotated images from 6 videos of 6 different infants aged from 17 weeks to 18 months. Two of these videos (infants) were recorded at home by the parents, with the infant sitting on the lap of a parent, themselves sitting on a couch. The remaining four videos (infants) were recorded in a laboratory context, with the camera placed in front of the participants, who were sitting (or standing) on a mat placed on the ground. The parent, and in particular their face, was not always fully visible in the images. The number of 359 images was reached after filtering down the images that were originally annotated but where the infant was missing (moved out of the field of the camera), or because an experimenter was in front of the camera and occluded most of the infant.
  • Real infants – SyRIP. Huang and colleagues created a hybrid dataset consisting of 1700 infant images: 1000 synthetic and 700 real infants, which they manually annotated and made public: the SyRIP dataset [55]. The subset dataset dedicated to benchmarking performance rather than training, which is what we used in our comparisons, contains 500 images of 50 infants up to 1 year old, in a large variety of contexts (at home, outside, in a laboratory), postures (supine, prone, standing, facing towards or away from the camera) and camera angles (see the original paper, which shows numerous examples). The images were taken from Youtube videos and Google Images results.
For real infants, it is difficult to estimate the difficulty of each sequence or image, though it is clear that our “lap” and the SyRIP datasets would be more difficult to estimate than our ”supine” dataset. The SyRIP dataset contains a wider variety of infant postures and camera angles, including many that partially hide the infants (e.g., from the side), making the estimation more difficult. As for the ”lap” dataset, since the methods were all originally trained on datasets of adult images, they are biased toward detecting adults, and the presence of the adult overlapping right behind the infant is a strong source of disturbance. For synthetic infants, the MINI-RGBD dataset has a definition of the sequence difficulty: easy (IDs 1 to 4), medium (5 to 9) and difficult (10 to 12), with this difficulty judged primarily with regard to the posture of the infant. Although the main manuscript focuses on the average results, it might be interesting to look at the results at the level of individual videos or their estimated difficulty. Such results are available in the Supplementary Materials.
A second coder annotated 20% of our manually-annotated image dataset. Intraclass Correlation Coefficients (ICC) were calculated for each keypoint and for each axis independently to measure the reliability of the human annotations using the Python module pingouin. The lowest ICC was 0.91 for the y-axis of the Left Hip keypoint, with the lower bound of its 95% Confidence Interval at 0.89. The mean ICC was $0.97\pm 0.03$, with all p-values $\ll 0.001$.
Our supplementary materials include the results for an extended version of our ”supine” dataset: 720 supplementary images from 8 additional videos; as well as a continuous sequence of 900 images (30 seconds) of a single infant at 17 weeks old. A summary table of each dataset’s characteristics is available in Sup. Tab. 1.
Comparison metrics
The metrics used for comparisons are described below. Because we observed differences in the results when the methods were given input in the form of videos or images, we present the results for both input types independently, except for DeepLabCut and ViTPose (within the MMPose environment) that only accepted video input.
The image-reference error that we compute as a building block for some of the metrics is the Euclidean distance between an estimated keypoint and its ground truth. Note that this ignores any gaps in the data when a method is not able to provide a keypoint’s estimated coordinates (see the Missing Data metric), or when ground truth is missing because the body part is occluded, as no distance can be computed in such cases.
Out of all these metrics, the ones that can be used even without ground truth, and for which we have the results on the full videos on our ”supine” dataset are: percentages of missing data and redundant detections, and the processing speed.
Object Keypoint Similarity (OKS)
Given $K$ keypoints that are both annotated for ground truth and detected by the pose estimation method, with $k\in[1,K]$: $d_{k}$ is the Euclidean distance between an estimated keypoint $p_{k}={(x_{kp},y_{kp})}$ and its ground truth $gt_{k}={(x_{gt},y_{gt})}$; $s$ is the area of the bounding-box around the target: a rectangle that encompasses all visible parts of the target within the image. As we do not have manually annotated bounding-boxes, we computed this area by taking the minimum and maximum X and Y values from all ground-truth keypoint coordinates in the image as an approximation. The coefficient $c_{k}$ is specific to each keypoint location and gives more weight to keypoints of which the ground-truth position varies less between human coders. We used the values recommended by the COCO challenge [56], available on its website, which are based on the annotations of the COCO dataset and give the most weight to the eyes and nose, and the least weight to the hips. From these, keypoint similarity $ks$ is computed for each estimated keypoint. OKS is the average $ks$ for a detection.
Average Precision (AP) and Average Recall (AR)
To evaluate the performance of pose estimation, the gold standard metrics in the literature are Average Precision and Average Recall [56], ranging between 0 and 100, higher values indicating better performance. We follow the benchmark evaluation set by the COCO challenge to calculate AP and AR, averaging the values computed over ten thresholds from 0.5 to 0.95, with steps of 0.05. Roughly, if the OKS for a detection is above the threshold and there is any ground truth that has not been matched with a detection yet on a given image, then the detection counts as True Positive. Other cases count as False Positives. If there is no detection but there is a ground truth, then it counts as a False Negative. Precision and Recall are computed for each threshold, then averaged. Note that despite the name Average Precision, it is actually computed as the area under the Precision-Recall Curve[57, 56]. In our case, the computation is simplified, as all our annotated images contain one single infant (one ground truth). For our ”Lap” dataset, the ground truth annotations of the parents were ignored, so that each method was only evaluated against the infant annotations.
Neck-MidHip error
As an alternative to OKS, we additionally use a reference distance, the Neck-MidHip length, corresponding roughly to the length of the infants’ torso. With this reference distance, we normalize the errors between recordings to compensate for different camera settings and infants of different ages and with different body sizes. Others and ourselves have previously used a similar strategy in a similar context [29, 11, 42]. Unlike OKS, such errors are easier to grasp as they directly relate to a measurable part of the body of the infant. Lower is better, and the minimum is 0.
The equations are shown in Eq. 2. Given $j$ a recording with $I_{j}$ annotated images for which Neck and MidHip keypoints ground truth are available or can be approximated. Given the images from the recording $i\in[0,I_{j}]$, then the Neck-MidHip length for this image $nmh\_len_{ji}$ is defined in pixels as the Euclidean distance between the Neck and the MidHip ground-truth coordinates in this image. For our ”supine” and the MINI-RGBD dataset, we take the median of all $nmh\_len_{ji}$ as the NeckMidHip length to use to normalize the recording: $nmh\_len_{j}$. Given a keypoint $k$ that is visible in the image and annotated, and detected by the pose estimation method, then its Neck-MidHip error $nmh\_error_{kij}$ is its error (Euclidean distance) relative to its ground truth, $d_{kij}$, divided by the median Neck-MidHip length for the recording, $nmh\_len_{j}$. For our ”lap” and the SyRIP dataset, which contain images from different sources mixed together, we normalize using each images’ neck-MidHip distance $nmh\_len_{ji}$ instead of a median from a whole recording. This has the downside that only images where this length can be estimated are used to calculate the error. Then, we obtain per-keypoint-location mean Neck-MidHip errors, and an overall mean full-body error, from all images and recordings.
$\begin{split}&nmh\_len_{ji}=\sqrt{(x_{neck}-x_{midhip})^{2}+(y_{neck}-y_{midhip})^{2}}\qquad nmh\_len_{j}=med(nmh_{ji})\\ &nmh\_error_{kij}=\frac{d_{kij}}{nmh\_len_{j}}\end{split}$ (2)
For the synthetic infants in the MINI-RGBD dataset, the Neck and MidHip positions are provided directly in the ground truth. In the real infants datasets, the ground truths do not provide Neck or MidHip keypoints. Instead, the centre of the right and left Shoulder and Hip keypoints are used, turning into a “MidShoulder-MidHip” length. We will simply write Neck-MidHip for simplification in the rest of the paper. The Neck-MidHip length is computed using the median of the Neck-MidHip lengths for a whole video when possible. This enables the use of all annotated images and keypoints even if one of the keypoints needed to compute the length is missing from the ground truth in a particular image (occlusions). It is also less subject to noise in the annotations, with the drawback that a single length is used for a whole video and does not fit exactly to each individual image. Note that using the median can only be done when there are no changes in camera settings, in particular zoom or camera position, which can happen, for example, when recordings are performed at home by a parent holding a smartphone with their hand; otherwise, it requires to separate the recording into subsets according to the different settings, or use per-image normalization instead, which is what we opted to do for our ”lap” and the SyRIP datasets.
Percentage of missing data
Missing data is especially undesirable for motion analysis, which requires accurate and consistent sampling over time. Missing data can happen in two ways. First, missing detections, when a method is not able to detect an infant in the image and does not provide any keypoint at all. Second, missing keypoints, when a method provides at least one detection, but fails to provide coordinates for one or more individual keypoints. An example is given in Fig. 1 (d), where the left foot keypoint is missing. This can happen either because of an erroneous estimation from the method or when body parts are occluded. Average Recall takes into account missing detections, but no other metric takes into account missing keypoints.
The total percentage of missing data computation is described in Eq. 3. We define the maximum amount of data $maxdata_{m}$ that can be estimated by a method $m$ as the number of images to estimate $N$ times the number of keypoints provided by the method $m_{kp}$. Missing detections $mdet$ are converted to this unit by multiplying with $m_{kp}$ as well, since any missing detection is equivalent to missing all keypoints in this one image. This number is added to the count of individual missing keypoints $mkp$, and divided by $maxdata_{m}$ to obtain the total percentage of missing data for this method: $mdata_{m}$.
$maxdata_{m}=N\times m_{kp}\qquad mdata_{m}=((mdet_{m}\times m_{kp})+mkp)/maxdata_{m}$ (3)
Percentage of redundant detections
Some methods can erroneously detect more people than there are in the image, as shown in Fig. 1 (b), where an additional ”person” is detected. These redundant detections might indicate that the method hallucinates other people on the image where there are none, or provides several estimations for the same person with different positions. Such behavior is undesirable, in particular for applications with more complex environments that, for example, involve interaction with an experimenter or the parents, such as in our ”lap” dataset, as it complexifies tracking each separate individual when ground truth is not available. The percentage of redundant detections is computed as the ratio between i) the difference in the number of detections between the expected and provided detection numbers (number of estimated detections minus the number of expected detections) divided by ii) the number of images the method provided detections for (total number of images minus the number of missed detections). For example, if there are 150 detections for 100 input images with 100 expected detections (1 infant per image), then 50 detections would be considered redundant, resulting in a percentage of 50% of redundant detections, translating that, on average, the method provides one additional detection every two images where it detected something. A percentage of 100% would mean that, on average, the method always provides two detections for each image, when one is expected. DeepLabCut, MediaPipe, and OpenPose did not have redundant detections (MediaPipe and OpenPose, have a built-in parameter to limit the detections to 1 person per image).
In the case of the ”Lap” dataset, where an adult is always present behind the infant, there are also cases where only the adult and not the infant is detected (Fig. 1 (f)). We empirically observed that some pose estimation methods have a tendency to often detect only the adult and not the infant in such context. This would show as a negative ratio of redundant detections. Taking the example from above, if a method had provided 150 detections for 100 images where 200 detections were expected due to the presence of both an infant and an adult, then we would get -50% “redundant” detections: on average, that method provided the two expected detections for half of the images, and only provided one detection for the other half. For this particular dataset, this is not a foolproof measure, as it is possible that a method detected the adult and then provided a second, redundant, detection of the adult or hallucinated another person somewhere else in the image instead of detecting the infant.
Percentages of redundant detections closer to 0 are preferred. Average Precision partially accounts for redundant detections: redundant detections with high scores modify AP values, but redundant detections with low scores do not.
Correlations between scores and OKS
In a real application scenario, without available ground truth, confidence or score values can be useful to provide cues about detections or keypoints that should not be trusted. For example, confidences are used as coefficients in the optimization process of smplify-x [58] when finding the parameters of the 3D full-body pose estimation from the 2D pose estimation. These values should be positively correlated with the OKS (the higher the OKS, the higher the score). When available, we use the score provided by the methods for each image. When not available, we used the median of the confidences as an alternative score for the whole detection. MediaPipe could not be evaluated as it does not provide scores or confidences. Using its implementation within the Python package scipy.stats, we measured correlations with Spearman Rank Order Coefficients.
Processing speed
Hardware and methods’ code modifications
All methods were run while turning off the display of visualizations on the screen and without saving the visualizations to the disk. To do so, some of the code of Detectron2 and the MMPose environment was modified, as no parameter was provided to turn them off. Further modifications were made to the Detectron2 code to output the keypoint coordinates as a file, as it did not provide a way to get such output by default. The ”lap” and SyRIP datasets contain non-sequential images extracted from different videos of different infants at different ages. The video input option was thus not available. On VITPose which in the version of MMPose we used could only process videos, we modified the code of MMPose’s ViTPose to accept images as inputs, without modifying any of the underlying processing done by ViTPose. While that enabled us to present ViTPose results on these datasets, ViTPose might internally only expect videos and have processing steps that enhance estimations assuming a time coherence between two processed frames. As there was not any parameter available to indicate to ViTPose a lack of sequential coherence, we avoided using the image processing steps for the other two datasets that had videos that could be processed (”supine”, and MINI-RGBD). The performance of ViTPose on the ”lap” and SyRIP dataset could be lower than would be in a real application scenario with full video recordings to process. All modifications are effective in the Dockers that we share (at https://hub.docker.com/u/humanoidsctu) and are indicated in the documentation.
@@ Results
The results for both real and synthetic infants are shown for each metric described above. In each subsection, we provide the results for the manually-annotated dataset of real infants first: in order, our ”supine” and ”lap” datasets, then the SyRIP dataset, and then the results for the MINI-RGBD dataset of synthetic infants are presented last. Due to the varying complexity of infant postures across the datasets, we provide detailed results for individual sequences in the Supplementary Materials when relevant.
Object Keypoint Similarity
The average OKS values for each dataset (average across individual infant images) are shown in Tab. 2 (see Supplementary Tables ST. 5 and ST. 6 to see the full details per video on the ”supine” and MINI-RGBD datasets).
For real infant images in ”supine” position, ViTPose and HRNet TD have the highest OKS, followed by HRNet BU. The results between frame-by-frame and video inputs are almost identical, except for OpenPose, for which image input leads to better results. On the ”lap” and SYRIP datasets, we observe a significant drop of performance for all methods, with particularly high variance, and wider gaps of performance between the methods. HRNet BU seems to perform the best on the ”lap” dataset, while ViTPose leads on the SyRIP dataset. For synthetic infants, VitPose and HRnet TD also show the best results. The differences between image and video inputs are narrow but consistent between the different methods, with a slight edge for image inputs. Synthetic infants in supine position are more difficult to process for all methods, in particular for HRNet BU, compared to the real infants in supine position, with the exceptions being DeepLabCut and MediaPipe, which perform better on synthetic infants, but the performance is overall higher than for the ”lap” or the SyRIP datasets.
ours – Sup. Videos 0.87$\pm$0.08 0.12$\pm$0.12 0.86$\pm$0.11 0.40$\pm$0.18 0.90$\pm$0.07 0.92$\pm$0.05 0.79$\pm$0.19 0.92$\pm$0.04
[…]
MINI-RGBD Videos 0.81$\pm$0.15 0.43$\pm$0.20 0.81$\pm$0.12 0.47$\pm$0.22 0.81$\pm$0.17 0.86$\pm$0.10 0.81$\pm$0.12 0.87$\pm$0.07
Average Precision and Average Recall
The details of the AP and AR values are shown in Tab. 3.
For real infants, ViTPose has the highest AP and AR. Considering only AP, ViTPose is followed by HRNet BU and then by OpenPose. Lower AP values for HRNet TD and Detectron2 in our supine dataset are influenced by a large number of redundant high-confidence detections (see Tab. 6). Considering only AR, ViTPose is followed by HRNet TD and HRNet BU. Whether the inputs are images (frames) or videos, the results are roughly similar except for OpenPose, though it seems that, generally, image inputs lead to better results (except for HRNet TD, for which video input seems preferred). There is a large drop of AP and AR in the ”Lap” and SyRIP datasets compared to the supine infants.
For synthetic infants, both the AP and AR values are lower than those in the supine dataset of real infants; it seems that the methods have more difficulty processing synthetic infants, except for DeepLabCut and MediaPipe. Considering only AP, ViTPose is the best method, followed by OpenPose. Considering only AR, HRNet TD is the best method, followed by ViTPose. We observe again that despite the lower scores, the synthetic infants of the MINI-RGBD datasets, which are also in supine position and seen from above, are better estimated by all methods than the real infants in the lap and in the SyRIP datasets.
ours–Supine images
AP 67.7 N/A 52.6 1.8 75.3 59.7 74.4 N/A
AR 74.7 N/A 79.4 6.1 85.1 89.3 79.0 N/A
ours–Supine videos
AP 67.5 0.0 24.1 1.8 74.6 60.5 55.2 88.5
AR 74.3 0.0 78.7 5.9 84.8 89.2 62.4 90.9
[…]
AP 25.0 N/A 11.0 7.8 10.9 20.1 27.0 30.4
AR 57.6 N/A 49.4 20.3 60.1 69.5 67.4 60.7
[…]
AP 28.0 N/A 25.6 0.0 26.2 30.9 38.4 41.9
AR 50.3 N/A 57.1 0.0 53.8 63.8 54.2 65.4
MINI-RGBD images
AP 60.5 N/A 48.1 6.1 62.2 59.9 66.5 N/A
AR 69.7 N/A 73.9 17.8 76.7 81.7 72.5 N/A
MINI-RGBD videos
AP 52.1 3.2 29.5 5.6 53.8 56.8 61.2 73.7
AR 62.7 12.3 68.0 17.4 71.6 77.9 67.5 79.1
@@
Neck-MidHip errors are shown in Fig. 2. DeepLabCut errors are not shown, as they create large radius circles that impede readability of the figures. The version of the figures including DeepLabCut is available in the Supplementary Figure SF. 2. Across all methods, the estimations of the eyes and nose are the most accurate, leading to the smallest errors, followed by shoulders and wrists. The positions of the hips and knees are detected with the highest errors. We can observe the increase in errors between the real infants in the ”supine” condition, the MINI-RGBD dataset, the SyRIP dataset, and finally the ”Lap” dataset where the errors make the image difficult to read. The details of the overall average Neck-MidHip errors across all keypoints with their standard deviations for each method are shown in Tab.  4. The variability of the errors is high, often above one third, and sometimes even half, of the average error. This might be explained by the large variability between each keypoint type as can be seen on Fig. 2, or Supplementary Materials Tab. ST 7, 8, 9 and 10.
Overall, for real infants, the best method is ViTPose, followed by HRNet TD. However, in the ”Lap” dataset, while the performance of all the methods drops significantly, ViTPose and the other top-down methods seem to struggle particularly compared to their performance on the other datasets. HRNet BU has the lowest errors among all methods, though the errors are high in absolute terms: an average error of 24% of the Neck-MidHip length. A less dramatic but general drop of performance is observed in the SyRIP dataset as well, where the best performing method is ViTPose with relatively high average errors at 12% of the Neck-MidHip length. For synthetic infants, the best method is HRnet TD, followed by ViTPose, although HRNet TD has slightly higher standard deviations.
Figure 2: Neck-midhip errors for each keypoint with available ground truth. The centre of the circles is the ground-truth position for that keypoint. The radius of each circle shows the average error as a proportion of the Neck-MidHip length (see Methods). The colors represent separately each pose estimation method. a) Real Infants ”supine”, image input; b) Real infants ”supine”, video input; c) MINI-RGBD synthetic infants, image input; d) MINI-RGBD synthetic infants, video input; d) Real infants ”lap”; and e) Real infants, SyRIP. DeepLabCut estimations leading to much bigger errors not shown (see Supplementary Fig. SF. 2).
@@
Real Images $8.3\pm 3.1$ N/A $10.2\pm 5.2$ 38.0$\pm$13.8 $7.8\pm 3.8$ 6.7$\pm$3.6 $8.8\pm 4.1$ N/A
Supine Videos $8.3\pm 2.9$ $99.6\pm 25.9$ $10.5\pm 5.4$ 35.1$\pm$13.7 $7.9\pm 3.9$ $6.7\pm 3.6$ $12.3\pm 4.3$ 6.0$\pm$2.7
[…]